OpenStudy (anonymous):

Solve for x using the addition principle: -8+x= -24 would that be -8x=-24 which equals 3?

OpenStudy (cwrw238):

you add 8 to each side of the equation -8 + x + 8 = -24 + 8

OpenStudy (cwrw238):

what happens to the lefthand side? - what does it simplify to?

OpenStudy (cwrw238):

-8 + 8 = ?

OpenStudy (anonymous):

16

OpenStudy (cwrw238):

no - 8 + 8 = 8 - 8 = 0

OpenStudy (cwrw238):

take the thermometer analogy: